OpenStudy (anonymous):

We can also use the equation for enthalpy change for physical phase changes. Consider the phase change H2O(l) → H2O(g). Calculate ΔHrxn. Use ΔHf values from the table above. ΔHrxn = A) -527.6 kJ B) 44.0 kJ C) 527.6 D) -44.0

OpenStudy (anonymous):

@touseii45 @thomaster

OpenStudy (anonymous):

where is the table?

OpenStudy (anonymous):

right here

OpenStudy (anonymous):

or this All the coefficients are 1. ΔHrxn = (sum of ΔHf products - sum of ΔHf reactants) ΔHrxn = [(1)(-1206.9 kJ/mole) + (1)(-241.8 kJ/mole)] - [(1)(-986.09 kJ/mole) + (1)(-393.6 kJ/mole)] We must be very careful in our math to properly handle all the negative signs. ΔHrxn = [-1206.9 + (-241.8)] - [-986.09 + (-393.6)] kJ ΔHrxn = [-1206.9 - 241.8] - [-986.09 - 393.6] kJ ΔHrxn = (-1448.7) - (-1379.69) kJ ΔHrxn = -1448.7 + 1379.69 kJ ΔHrxn = - 69.0 kJ This reaction is exothermic.

OpenStudy (anonymous):

uhm okay wow i understood none of that, @electrokid can you help?

OpenStudy (anonymous):

or @thomaster can you help? your a genius...

OpenStudy (anonymous):

i have no idea. here, help with this instead http://openstudy.com/study#/updates/51658bfde4b066fca66102c3

OpenStudy (anonymous):

Hint: remember that this will jsut be your Latent hear of vaporization ...

OpenStudy (anonymous):

just a phase change.. no change in temperature.

OpenStudy (anonymous):

\[285.8-241.8\]

OpenStudy (anonymous):

http://openstudy.com/study#/updates/51659543e4b0902dbc1d05b2 i really need help here